OpenStudy (anonymous):

help? Seats in an auditorium are numbered 1-500, and the principal is handing out snacks in an assembly. Every student whose seat is a multiple of 4 gets chocolate. Every student in a seat that is a multiple of 7 gets jelly beans, and every student in a seat that is a multiple of 10 gets chips. The student in which seat number is the first to get all 3 treats? -40 -70 -100 -140 what should i do to get the answer?

Directrix (directrix):

I think we want the LCM (least common multiple) of 4, 7, and 10.

Directrix (directrix):

LCM The smallest positive integer into which two or more integers divide evenly. For example, 24 is the LCM of 8 and 12. Sometimes the LCM is written using brackets: [8, 12] = 24.

Directrix (directrix):

What is the smallest positive integer into which 4, 7, and 10 will divide?

Directrix (directrix):

4 = 2*2 7 = 7*1 10 = 2*5 To get the LCM, you need the product of 2, 2, 7, and 5. What is that number? @alynck

OpenStudy (anonymous):

140?

Directrix (directrix):

@alynck That's what I got.